OpenStudy (anonymous):

as I can do this? Expand as a sum or diference: long3 (xy)3

OpenStudy (triciaal):

do you mean log base 3 of (xy)^3 or something else?

OpenStudy (triciaal):

can you clarify the question?

OpenStudy (triciaal):

sorry can't help if I don't know what you are asking and you are not responding.

Nnesha (nnesha):

its \[\large\log _{3}(xy)^{3}\]

OpenStudy (triciaal):

@Nnesha that's what I asked

Nnesha (nnesha):

no i'm saying yes that is question 3 is exponent

OpenStudy (jhannybean):

well, remember. \[\log(a) +\log(b) =\log(ab)\]

OpenStudy (jhannybean):

\[\log\left(\frac{a}{b}\right) =\log(a) - \log(b)\]

OpenStudy (anonymous):

\[\Large \log(a^{p})=p \times \log(a)\]

OpenStudy (triciaal):

logs are easy but I don't appreciate the asker not responding

OpenStudy (jhannybean):

That's true, :)